Can you solve this by graphing?
y = -3x+3
6x+y = 6
The solution set is what? Type an ordered pair ( )

Thank you!

Answer by mananth(16949)

-3 x -1 y = -3................(1)
6 x 1 y = 6.....................2)
equation1 3 = 3
when x= 0, y= 3, ( 0 , 3 )
when y = 0, x= 1, ( 1 , 0 )
Equation 2
when x= 0, y= 6, ( 0 , 6 )
when y = 0, x= 1, ( 1 , 0 )


Solution (1,0)